# CMoE: Fast Carving of Mixture-of-Experts for Efficient LLM Inference

### v1

@@ V-B Ablation Studies
TABLE II: Comparison of downstream tasks. Randomly select samples from WikiText-2 training dataset as calibration and training data.
@@
Model Method Type FT BoolQ(32) SciQ PIQA WinoGrande(5) ARC-C(25) HellaSwag(10) LLaMA-2-7B Dense - - 82.04 90.80 78.78 73.95 53.15 78.55 LLama-MoE A2E8 37.83 20.00 49.73 50.12 25.79 26.18 LLama-MoE A4E16 37.82 20.60 49.56 49.40 25.98 27.21 CMoE S1A1E8 46.09 65.30 52.77 48.70 23.80 30.12 CMoE S1A3E16 50.79 62.00 52.29 50.12 24.23 28.76 CMoE S2A2E16 53.85 66.50 53.26 49.33 23.72 29.85 CMoE S1A1E8 55.04 77.50 57.12 54.06 27.56 38.79 CMoE S1A3E16 53.66 75.30 56.37 53.98 27.82 38.89 CMoE S2A2E16 57.19 77.30 56.86 52.57 26.45 38.90 LLaMA-3-8B Dense - - 83.48 94.2 80.79 77.50 59.72 82.16 LLama-MoE A2E8 37.82 20.30 49.02 50.74 25.68 25.76 LLama-MoE A4E16 37.83 20.50 51.52 49.96 25.27 26.00 CMoE S1A1E8 47.43 47.70 52.56 50.67 23.72 28.08 CMoE S1A3E16 41.31 46.60 51.63 48.77 24.15 27.41 CMoE S2A2E16 45.08 53.10 50.87 53.19 24.23 28.44 CMoE S1A1E8 56.88 72.00 57.34 51.93 25.77 36.68 CMoE S1A3E16 60.55 71.40 57.07 53.51 26.54 36.28 CMoE S2A2E16 62.14 72.20 59.08 51.14 27.73 36.68
Figure 5: Ablation studies:(a) Impact of shared expert ratio on model performance; (b) Activation Rate vs. Model Performance; (c) Effect of Load Balancing.
Impact of Training Data Size. To quantify the relationship between training data volume, model performance, and computational efficiency, we systematically vary the number of training samples from 0 (initial state) to 4,096 while measuring perplexity (PPL) and construction time. Fig. 4 illustrates the results on the setting of “S2A2E16”, revealing critical trends in performance-cost trade-offs.
Increasing training data from 0 to 4,096 samples reduces perplexity by 80.4% (62.30 $\to$ 12.21) but non-linearly increases construction time from 298s to 4,502s. While performance improves sharply initially (64 samples cut PPL by 63%), gains diminish beyond 1,024 samples (13.47 PPL), with marginal improvements (12.21 PPL at 4,096 samples) requiring disproportionately higher runtime (+133% from 2,048 to 4,096 samples). The results demonstrate that CMoE only needs a modest of data to achieve low perplexity, while further performance increase is hard and possibly demands more diverse data.
Shared Experts Ratio. We carry out an extensive analysis of the influence of the shared expert ratio on perplexity. In our experimental setup, we have a total of 32 experts, among which 8 are activated. We systematically adjust the proportion of shared experts within these 8 activated experts. As depicted in Table Fig. 5, the results clearly demonstrate a distinct trend: as the ratio of shared experts rises from 0.125 to 0.875, the perplexity continuously decreases. There is a remarkable improvement in the perplexity value, dropping from 14.48 to 11.93. This finding indicates that shared experts generally contribute positively to the performance of the model. However, it is also evident that the marginal returns tend to diminish when the ratio of shared experts reaches a relatively high level.
Activation Rate. We conduct an external analysis to investigate how the total activation rate (comprising both shared and activated routed experts) impacts model performance across different domains, in comparison with dense baselines. We perform evaluations on the CMoE model featuring a total of 16 experts. Specifically, we vary the ratio of shared experts and activated routed experts at a fixed proportion of 1:1 on the WikiText-2 and C4 datasets. As illustrated in Table Fig. 5, a monotonic decrease in PPL can be found as the activation rate increases, gradually approaching the performance level of the dense model. For both datasets, an activation rate of 75% enables the model to achieve nearly comparable performance to the dense model. On the WikiText-2 dataset, the PPL values are 5.79 for the model under consideration and 5.27 for the dense model; on the C4 dataset, the corresponding values are 11.19 and 7.27. These results demonstrate that CMoE architectures can attain a performance quality comparable to that of dense models, even when characterized by relatively high sparsity.
Load Balancing. CMoE can inherently achieve effective load-balancing among routed experts, with the exception of some special blocks. For example, the last block of LLaMA-2-7B exhibits extremely unbalanced expert counts. However, this issue can be effectively resolved by the load-balancing mechanism we introduced in Section IV-C. As presented in Fig. 5, prior to the implementation of the load-balancing strategy, substantial disparities in the workloads assigned to different experts are clearly observable. Specifically, Expert 3 and Expert 8 are burdened with disproportionately high loads, handling 11,163 and 11,225 instances respectively. In stark contrast, other experts process a significantly lower number of instances. For example, Expert 12 and Expert 13 handle only 100 and 102 instances respectively. Upon the application of our proposed load - balancing mechanism, a more equitable and uniform distribution of the computational workload across all experts is successfully achieved. After the adjustment, the number of instances processed by each expert falls within the range of 1,443 to 3,584.
VI Conclusion
@@
We present CMoE, a framework that efficiently carves sparse Mixture-of-Experts (MoE) architectures from dense LLMs through parameter reorganization and lightweight adaptation. By leveraging activation sparsity patterns in FFN layers, CMoE groups neurons into shared and routed experts via a novel balanced linear assignment formulation. The router is analytically initialized from activation statistics and refined via differentiable scaling and load balancing. Experiments show that CMoE can construct effective Mixture-of-Experts (MoE) models, achieving comparable perplexity to dense models and outperforming baseline models. For instance, on the SciQ dataset, CMoE can reach 56.3% even without fine-tuning and it can further improve the accuracy to 76.59% after fine-tuning, a huge outperforming compared with 21.2% (the accuracy of the baseline LLaMA-MoE). Extensive experiments have demonstrated that CMoE offers a practical approach for deploying large language models (LLMs) in resource-constrained environments.

### v2

@@ 4.1 Main Results
Table 2: Comparison of downstream tasks. Higher is better. Randomly select samples from WikiText-2 training dataset as calibration and training data.
@@
Model Method Act. FT BoolQ(32) SciQ PIQA WinoGrande(5) ARC-C(25) HellaSwag(10) LLaMA-2-7B Dense - - 82.04 90.80 78.78 73.95 53.15 78.55 LLama-MoE 25% 37.83 20.00 49.73 50.12 25.79 26.18 LLama-MoE-v2 25% 51.25 67.00 56.64 52.88 25.68 35.10 CMoE 25% 46.09 65.30 52.77 48.70 23.80 30.12 CMoE 25% 55.04 77.50 57.12 54.06 27.56 38.79 LLama-MoE 75% 37.83 20.00 49.35 50.28 26.37 25.77 LLama-MoE-v2 75% 49.24 79.50 63.55 59.35 34.81 54.89 CMoE 75% 71.68 88.60 73.94 65.90 41.98 69.21 CMoE 75% 75.72 89.60 74.34 67.40 43.94 70.36 LLaMA-3-8B Dense - - 83.48 94.2 80.79 77.50 59.72 82.16 LLama-MoE 25% 37.82 20.30 49.02 50.74 25.68 25.76 LLama-MoE-v2 25% 48.53 45.80 50.87 48.38 24.91 27.31 CMoE 25% 47.43 47.70 52.56 50.67 23.72 28.08 CMoE 25% 56.88 72.00 57.34 51.93 25.77 36.68 LLama-MoE 75% 37.83 20.60 51.31 49.80 26.19 26.08 LLama-MoE-v2 75% 53.30 50.00 50.00 51.38 24.49 28.42 CMoE 75% 58.29 89.90 72.36 65.03 47.52 66.94 CMoE 75% 69.20 88.70 72.80 66.30 48.80 70.63
Evaluation on PPL. Table 1 presents the perplexity values for the WikiText-2 merity2016pointer and C4 raffel2020exploring validation sets. In the training-free scenario, CMoE achieves approximately 60 on WikiText-2 when 25% of its components are activated, thereby showcasing its immediate practical utility, but LLaMA-MoE and LLaMA-MoE-v2 both suffer from catastrophic degradation, resulting in perplexity values greater than 20,000 or even NaN without any training. When subjected to a lightweight 2k fine-tuning process, CMoE experiences a notable reduction in perplexity across all activation ratios as expected. Specifically, 75% activation setting allows CMoE to perform on par with the dense model, evidenced by nearly PPL values. Further enhancements through extended fine-tuning is explored in the Ablation Studies section.
Evaluation on Downstream Task Performance. Table 2 presents zero-shot and few-shot results on benchmarks including: 32-shot BoolQ clark2019boolq , 0-shot PIQA bisk2020piqa , 0-shot SciQ welbl2017crowdsourcing , 5-shot Winogrande sakaguchi2021winogrande , 25-shot ARC-Challenge clark2018think , and 10-shot HellaSwag zellers2019hellaswag . Similar trends emerge: CMoE is functional training-free, substantially outperforming other methods like LLaMA-MoE and LLaMA-MoE-v2 that collapse without training. Lightweight 2k fine-tuning yields significant gains; for the $25\%$ activation CMoE, we recover $>76\%$ of the dense model’s accuracy on SciQ and BoolQ. Increasing the activation ratio further closes the gap to the dense model. Comparisons with Dense-Small models and the impact of extended fine-tuning are detailed in the Ablation Studies (Section 4.2).
Evaluation on Inference Speedup. We measure inference latency on one single GPU. Table 3 shows speedups for the MLP layers and the full model end-to-end. Even with the overhead of routing and non-MLP layers, CMoE achieves substantial end-to-end speedups, ranging from 1.4×1.4$\times$1.4 × to 1.6×1.6$\times$1.6 × depending on batch size and configuration. This validates that CMoE can achieve practical efficiency enhancements, even though there is a slight performance compromise at high levels of sparsity.
Table 3: Measured inference speedup for various CMoE activation configurations. S: Shared experts; A: Active routed experts; E: Total experts. E.g. S1A1E8 (1 shared + 1 active routed) / 8 total.
Configuration Act. Memory-Bound Compute-Bound FFN Speedup Full Model Speedup FFN Speedup Full Model Speedup S1A1E8 $25\%$ 3.6×3.6$\times$3.6 × 1.5×1.5$\times$1.5 × 3.75×3.75$\times$3.75 × 1.6×1.6$\times$1.6 × S1A3E16 $25\%$ 3.2×3.2$\times$3.2 × 1.4×1.4$\times$1.4 × 3.45×3.45$\times$3.45 × 1.5×1.5$\times$1.5 × S3A3E8 $75\%$ 1.25×1.25$\times$1.25 × 1.05×1.05$\times$1.05 × 1.27×1.27$\times$1.27 × 1.12×1.12$\times$1.12 × S3A9E16 $75\%$ 1.15×1.15$\times$1.15 × 1.02×1.02$\times$1.02 × 1.09×1.09$\times$1.09 × 1.05×1.05$\times$1.05 ×
4.2 Ablation Studies
We conduct ablations using Llama-2 7B and WikiText-2 data unless otherwise noted.
Impact of Fine-tuning Data Size. Figure 2 illustrates the trade-off between fine-tuning data size (using WikiText-2 samples) and construction time versus perplexity (PPL) for the CMoE $25\%$ activation configuration. Performance improves rapidly initially when increasing data from 0 to 1,024 samples, but gains diminish significantly afterwards, while the time cost increases non-linearly. For instance, increasing samples from 2,048 to 4,096 only marginally improves PPL (12.72 to 12.21) but increases runtime by 133%. This suggests that lightweight fine-tuning (e.g., 2k samples) is highly effective, and substantial further gains likely require more diverse datasets like SlimPajama (see Section 4.2) rather than simply more of the same fine-tuning data.
Figure 2: Trade-off between Model Performance (PPL) and Construction Time with Increasing Fine-tuning Data (WikiText-2 samples, $25\%$ activation).
Load Balancing. Figure 3 demonstrates the effectiveness of our load balancing mechanism (Section 3.3) on the final layer of Llama-2 7B, which exhibits highly skewed activation without balancing. The mechanism successfully distributes load more evenly across experts.
Figure 3: Effect of Load Balancing on expert utilization in Llama-2 7B final block ($25\%$ activation).
Comparison with Dense-Small Models and Extended Fine-tuning. Table 4 compares CMoE ($25\%$ activation) against Pythia-1.0B and TinyLlama-1.1B, which have roughly similar active parameter counts. The table also shows CMoE performance after fine-tuning on up to 1.2B tokens from the diverse SlimPajama dataset. Training-free CMoE underperforms the pre-trained dense-small models. However, after lightweight 2k fine-tuning, CMoE becomes competitive. With extended fine-tuning (1.2B tokens), CMoE surpasses Pythia-1.0B on several benchmarks and approaches TinyLlama-1.1B, despite using vastly less training data than either (1.2B vs 299B/3T tokens). This demonstrates that CMoE effectively transfers knowledge from the base model and benefits significantly from further, albeit still relatively modest, fine-tuning.
Table 4: Comparison with smaller dense models after fine-tuning on SlimPajama. CMoE uses $25\%$ activation (S1A1E8 configuration).
Model Training Data PIQA BoolQ HellaSwag Winogrande Pythia-1.0B 299B 69.21 57.83 47.16 53.43 TinyLlama-1.1B 3T 73.29 57.83 59.20 59.12 CMoE-25% training-free 52.77 46.09 28.08 50.67 CMoE-25% 2k WikiText-2 57.12 55.04 36.68 51.93 CMoE-25% 0.5B SlimPajama 59.41 57.13 41.84 53.67 CMoE-25% 1.2B SlimPajama 61.64 59.51 44.02 55.88
Impact of Expert Configuration ($25\%$ Activation). We compare different expert configurations while keeping the total activation ratio fixed at $25\%$. Table 5 shows that configurations like S1A1E8 and S2A2E16, which balance shared and routed experts, generally perform well. S1A1E8 slightly edges out S2A2E16 in PPL, potentially due to simpler routing, while S2A2E16 shows slightly better performance on BoolQ. S1A3E16, with more complex routing (3 active routed experts), shows slightly worse performance.
Table 5: Performance of different CMoE configurations at a fixed $25\%$ activation ratio (Llama-2 7B).
Configuration WikiText-2 $\downarrow$ C4 $\downarrow$ BoolQ $\uparrow$ SciQ $\uparrow$ TF FT TF FT TF FT TF FT S1A1E8 60.86 12.76 135.61 32.12 46.09 55.04 65.30 77.50 S2A2E16 62.30 12.73 136.12 32.37 53.85 57.19 66.50 77.30 S1A3E16 89.19 13.84 180.65 33.56 50.79 53.66 62.00 75.30
5 Conclusion
@@
We present CMoE, a framework that efficiently carves sparse Mixture-of-Experts (MoE) architectures from dense LLMs through parameter reorganization and lightweight adaptation. By leveraging activation sparsity patterns in FFN layers, CMoE groups neurons into shared and routed experts via a novel balanced linear assignment formulation. The router is analytically initialized from activation statistics and refined via differentiable scaling and load balancing. Experiments demonstrate that increasing the activation ratio to 75% achieves remarkable results, delivering lossless precision from the perpective of the perplexity while still maintaining a 5% acceleration. After a brief LoRA fine-tuning process (requiring only 1 hour and 2,000 samples) successfully recovers over 76% of the dense model’s downstream accuracy, highlighting the practical efficiency of this approach for resource-constrained deployment scenarios, which outperforms the baselines LLaMA-MoE and LLaMA-MoE-v2. Extensive experiments have demonstrated that CMoE offers a practical approach for deploying large language models (LLMs) in resource-constrained environments.
Appendix A Algorithmic Analysis
Figure A.2: The proposed framework, CMoE, and the numerical analysis on its effectiveness.
As demonstrated in LABEL:fig:hs, the distribution of the FFN hidden state $\mathbf{h}$ is sharply peaked at 00 and constrained within a small range. This indicates that most $h_{i}$ are concentrated near zero, confirming the sparsity of activations. The symmetry of the distribution suggests a near-normal distribution centered around the mean. These observations of high differentiation and independence across different neurons suggest that hidden state values serve as a robust basis for judging neuron activation, motivating the Absolute Top-K (ATopK) metric defined in Equation 4.
To quantify neuron activation patterns, we first collect activation markers from a calibration dataset. Given a batched input tensor $\mathbf{X}\in\mathbb{R}^{b\times s\times d}$, with $b$ the batch size and $s$ the sequence length, we obtain the batched hidden state $\mathbf{H}\in\mathbb{R}^{b\times s\times d_{h}}$ as follows:
Note that in practical implementation, we normalize $\mathbf{X}$, $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$ before the calculation to eliminate the influence of their magnitudes on the output. Using the ATopK metric (defined in the main text), we then calculate the activation markers for the hidden state of all the tokens. These markers are reshaped into an activation feature matrix $\mathbf{A}\in\mathbb{R}^{q\times d_{h}}$, where $q=b\cdot s$. Denote the $i$-th column of $\mathbf{A}$ as the feature vector $\mathbf{c}_{i}\in\mathbb{R}^{q}$, i.e., $\mathbf{A}=[\mathbf{c}_{1}\;\mathbf{c}_{2}\;\cdots\;\mathbf{c}_{d_{h}}]$, which represents the $i$-th neuron’s activation status on the sampled $q$ tokens. By calculating the expected value of each $\mathbf{c}_{i}$, we can obtain the vector of activated rates $\boldsymbol{\mu}$ as:
$\boldsymbol{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}],\mbox{ with }\mu_{i}=\frac{1}{q}\sum_{j=1}^{q}c_{j,i},$ (28)
where $c_{j,i}$ indicates the $j$-th element of $\mathbf{c}_{i}$. The histogram of these activation rates $\boldsymbol{\mu}$ is shown in LABEL:fig:rates. The histogram reveals a highly skewed distribution of activation rates, where the majority of neurons exhibit low activation rates (below 0.1), with a sharp peak near 0.07. However, the distribution also features a long tail, indicating the presence of a subset of neurons with significantly higher activation rates extending up to 1. These high-activation neurons are likely active across a wide range of input tokens, making them suitable for processing common knowledge rather than task-specific specialization. Therefore, we identify neurons for shared experts by grouping these high-activation neurons. Given the total number of shared experts as $N_{s}$ and the expert size $m$, we get the selection indices set $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with highest activation rates based on $\boldsymbol{\mu}$:
These indices $S_{N_{s}}$ are then used to form the shared experts by assigning the corresponding parameters from the original FFN, as detailed in Section 3.1 (see Equation LABEL:eq:expert_w in the main text).
The majority of low activation rates also encourage us to construct routed experts, which are not always activated but are specialized for tokens encountered.
Appendix B Impact of Calibration Data Size
We vary the number of calibration samples (WikiText-2) used for neuron analysis. Table B.1 shows that performance improves modestly with more calibration data, with diminishing returns beyond 32 or 64 samples. Using 8 samples provides a good balance of efficiency and effectiveness.
Table B.1: Impact of calibration data size on PPL (Llama-2 7B, CMoE $25\%$ activation, training-free).
Calibration Samples WikiText-2 PPL C4 PPL
8 60.86 135.61
32 58.71 129.62
64 55.88 125.02
Appendix C Scaling to Larger Models
To verify that CMoE scales to larger models, we applied the same conversion process (using $25\%$ activation, S1A1E8 configuration) to Llama-2 13B. Table C.1 shows the training-free PPL results using different calibration datasets. We observe similar patterns to the 7B model: performance improves with more calibration data, and using domain-matched calibration data (C4 for C4 PPL) yields better results. This confirms the applicability of CMoE to larger base models.
Table C.1: CMoE ($25\%$ activation, training-free) performance on Llama-2 13B with varying calibration data.
Calibration Data WikiText-2 PPL C4 PPL
8 from WikiText-2 68.82 193.66
32 from WikiText-2 56.25 167.25
32 from C4 116.77 134.93
32 WikiText-2 + 32 C4 68.59 140.17
Appendix D Discussion
Limitations While CMoE demonstrates promising results in transforming dense LLMs into efficient sparse MoE architectures, certain aspects merit consideration for future exploration. Firstly, the performance of neuron activation profiling and subsequent expert grouping is influenced by the choice of the calibration dataset. Optimal results are best achieved when the calibration data aligns well with the characteristics of the target domain or downstream tasks. Secondly, while the training-free router construction provides notable computational efficiency, it may not fully capture all nuances of routing decisions in the same way a router trained extensively end-to-end with the experts might. In some scenarios, particularly those requiring highly dynamic expert selection, there might be opportunities for further performance gains through additional fine-tuning. Future work could also explore the application and adaptation of CMoE across a wider array of model architectures and a more diverse set of downstream tasks to further understand its broader applicability and potential refinements.
Impact Statement This paper aims to contribute to the advancement of the Machine Learning field. While our work may have various societal implications, we do not find it necessary to emphasize any particular consequences here.